while fishing, you notice that eight crests pass a given point in a time interval of 10s. You approximate the distance between two crests to be 1.2m. Calculate the


a. Speed of water waves and


b. Depth of water

1
Expert's answer
2022-01-24T11:09:32-0500

Given:

"t=10\:\rm s"

"l=8\lambda"

"\lambda=1.2\:\rm m"


(a) The speed of water waves 

"v=\frac{l}{t}=\rm\frac{8*1.2\: m}{10\: s}=0.96\: m/s"

(b) N/A